x+2y+3z=6
y+2z=0
z=2

Answer by stanbon(50455)

x+2y+3z=6
y+2z=0
z=2
================================
Solutions:
z = 2
-------
So, y + 2*2 = 0 ; So, y = -4
-------
Then, x + 2*-4 + 3*2 = 6
x - 8 + 6 = 6
x = 8
====================
Cheers,
Stan H.